n January 2nd, Jenny purchased a delivery van for cash. Jenny estimates that the business will use the van for 8 years and that it will have a resale value after 8 years for 30% of the purchase price. In addition to the purchase price of the van, Jenny had to pay $1,000 for tax and delivery charges. whats the depreciation expense using double declining method. The purchase price for the van is $21900.
Answer #1

Solution:

Cost of VAN = Purchase price + Tax and delivery charges = $21,900 + $1,000 = $22,900

Depreciation rate - SLM = 1/8 = 12.50%

Depreciation rate - Double declining = 12.50%*2 = 25%

Depreciation expense using double declining method for the year of purchase = $22,900*25% = $5,725